OpenStudy (anonymous):

What is the value of a? 5(a-4)-8a=55 a=?

OpenStudy (igreen):

5(a - 4) - 8a = 55 Distribute 5 into the parenthesis: 5a - 20 - 8a = 55 Simplify: -3a - 20 = 55 Add 20 to both sides: -3a = 75 Divide -3 to both sides: a = -25

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Samus_Aran

OpenStudy (anonymous):

Thank you